Suppose you have 25 years until you retire, and that you desire a retirement nest-egg of $2,500,000 on the day you retire. Suppose also that you’ve saved $100,000 toward your retirement so far, and that your investment account earns a nominal rate of 7.5% per year, compounded monthly. In addition, suppose you expect a windfall inheritance of $200,000 five years from now that you will invest in this account.

a) What is the effective interest rate, or annual percentage yield, of your investment account?

b) How much money should you deposit into that account every year in order to reach your goal?
Answer #1

1.
=(1+7.5%/12)^12-1=7.7633%

2.
=(2500000/(1+7.7633%)^25-100000-200000/(1+7.7633%)^5)*7.7633%/(1-1/(1+7.7633%)^25)=13586.0757